OpenStudy (anonymous):

Which expression is a fourth root of -1+isqrt3? a) fourth root of 2(cos(280degrees)+isin(280degrees)) B) fourth root of 2 (cos(30degrees)+isin(30degrees)) C) fourth root of 2 (cos60degrees)+isin(60degrees) D) fourth root of 2( cos320degrees)+isin(320degrees))

OpenStudy (dumbcow):

\[(-1 + i \sqrt{3})^{1/4} = [2(\cos 120 + i \sin 120)]^{1/4}\] \[ = \sqrt[4]{2} ( \cos \frac{120 +360k}{4} + i \sin \frac{120+360k}{4})\] \[k = 0,1,2,3\] ---> plug in each k value to get 4 roots Only one will match given answers